### v1

@@ 5 Related Work
Sequential recommender systems (SRS) [27] have become ever-increasing prevalent by learning precise user and item embeddings from sequential behaviors. Formally, numerous efforts have been devoted to designing sophisticated models for the SRS scenario. A well-known traditional work is the FPMC model, which combines the Markov chain and matrix factorization techniques for the next basket recommendation. Later, researchers largely facilitate this area with various advanced neural model techniques. As one pioneering attempt, GRU4Rec has been widely used to solve sequential recommendations by employing the hidden state of recurrent neural networks (RNNs) to propagate user preference. Due to its effectiveness, a series of extensions [15] have been proposed based on such models. In addition, self-attentive and convolutional-based models can be very effective in modeling sequential behaviors due to their strengths in the capacity of long-term dependence modeling and easy parallel computation. Recently, relying on historical session behavior, we notice that graph neural networks [17] have become increasingly prevalent by constructing graph structures from session sequences.

### v2

@@ 5 Related Work
Sequential recommender systems (SRS) [27] have become ever-increasing prevalent by learning precise user and item embeddings from sequential behaviors. Formally, numerous efforts have been devoted to designing sophisticated models for the SRS scenario. A well-known traditional work is the FPMC model, which combines the Markov chain and matrix factorization techniques for the next basket recommendation. Later, researchers largely facilitate this area with various advanced neural model techniques. As one pioneering attempt, GRU4Rec has been widely used to solve sequential recommendations by employing the hidden state of recurrent neural networks (RNNs) to propagate user preference. Due to its effectiveness, a series of extensions [16] have been proposed based on such models. In addition, self-attentive and convolutional-based models can be very effective in modeling sequential behaviors due to their strengths in the capacity of long-term dependence modeling and easy parallel computation.
@@ 6 Conclusion
Acknowledgement This research was supported by grants from the National Key Research and Development Program of China (Grant No. 2021YFF0901003) and the Fundamental Research Funds for the Central Universities.